# AmpleGCG: Learning a Universal and Transferable Generative Model of Adversarial Suffixes for Jailbreaking Both Open and Closed LLMs

### v1

@@ 1 Introduction
With large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023) demonstrating impressive performance on a wide range of tasks, more and more of them are being integrated into autonomous systems and deployed to solve real-world problems. Hence, improving their safety and trustworthiness has attracted significant attention (Yang et al., 2024; Tang et al., 2024; Mo et al., 2024) with a large amount of efforts (Xu et al., 2024; Yuan et al., 2024; Bhatt et al., 2023) focusing on developing safety guardrails.
@@ 4.2 Experimental Setup
Baselines. We compare our AmpleGCG with the GCG default setting under both the individual query and mutliple queries settings and its augmented counterpart. According to (Mazeika et al., 2024), AutoDAN (Liu et al., 2023c) serves as the second most effective approach to jailbreaking after GCG, we report the performance of AutoDAN-GA as well.
Target Models. We present the effectiveness of different attack methods in jailbreaking Vicuna-7B (Chiang et al., 2023) and Llama-2-7B-Chat (Touvron et al., 2023), given any harmful test queries. To show AmpleGCG’s transferability, we further include open-source Mistral-7B-Instruct (Jiang et al., 2023) and closed-source GPT-3.5-0613, GPT-3.5-0125, GPT-4-0613 models (OpenAI et al., 2023) as our target models.
@@ 4.3 Results
More excitingly, we delve into whether AmpleGCG, trained on open-source models, can be applied to attacking closed-source models. We collect training pairs from four models—Vicuna-7B, Vicuna-13B, Guanaco-7B, and Guanaco-13B (Dettmers et al., 2023), following the findings in (Zou et al., 2023). For details on our data collection methodology when optimizing across multiple models, please refer to Appendix B. For API cost consideration, we randomly select 50 queries from the 100 queries test set. We further incorporate two extra stringent evaluators for detecting harmfulness to reduce false positives: classifier from (Mazeika et al., 2024) and GPT-4 evaluator (evaluation prompt is in Appendix O). We take the GCG mutliple queries setting as our baseline, which is optimized using the aforementioned four models as well.

### v2

@@ 1 Introduction
With large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023) demonstrating impressive performance on a wide range of tasks, more and more of them are being integrated into autonomous systems and deployed to solve real-world problems. Hence, improving their safety and trustworthiness has attracted significant attention (Yang et al., 2024; Tang et al., 2024; Mo et al., 2024) with a large amount of efforts (Xu et al., 2024; Yuan et al., 2024; Bhatt et al., 2023) focusing on developing safety guardrails.
@@ 4.2 Experimental Setup
Baselines. We compare our AmpleGCG with the GCG default setting under both the individual query and mutliple queries settings and its augmented counterpart. According to (Mazeika et al., 2024), AutoDAN (Liu et al., 2023b) serves as the second most effective approach to jailbreaking after GCG, we report the performance of AutoDAN-GA as well.
Target Models. We present the effectiveness of different attack methods in jailbreaking Vicuna-7B (Chiang et al., 2023) and Llama-2-7B-Chat (Touvron et al., 2023), given any harmful test queries. To show AmpleGCG’s transferability, we further include open-source Mistral-7B-Instruct (Jiang et al., 2023) and closed-source GPT-3.5-0613, GPT-3.5-0125, GPT-4-0613 models (Achiam et al., 2023) as our target models.
@@ 4.3 Results
More excitingly, we delve into whether AmpleGCG, trained on open-source models, can be applied to attacking closed-source models. We collect training pairs from four models—Vicuna-7B, Vicuna-13B, Guanaco-7B, and Guanaco-13B (Dettmers et al., 2024), following the findings in (Zou et al., 2023). For details on our data collection methodology when optimizing across multiple models, please refer to Appendix B. For API cost consideration, we randomly select 50 queries from the 100 queries test set. We further incorporate two extra stringent evaluators for detecting harmfulness to reduce false positives: classifier from (Mazeika et al., 2024) and GPT-4 evaluator (evaluation prompt is in Appendix O). We take the GCG mutliple queries setting as our baseline, which is optimized using the aforementioned four models as well.